Home‎ > ‎Reports‎ > ‎Crew 108‎ > ‎

day13


Crew #108 Log Book for December 15, 2011

Photos of the day





Commander's Check-In Report

Date and time: 

15 December 2011

Written by : 

Charlotte Poupon, Crew 108, Commander

Crew Physical Status:

All 5 all right

Time Departed/Returned from EVA:

EVA # 22 from 10:11 to 12:46
EVA # 23 from 13:45 to 15:55

Brief Narrative of Field Mission Results:

Not applicable

EVA Data/Interpretations:

Not applicable

Engineering/Hab Maintenance:

Engineer to fix the breathing gear damaged two days ago.

Report Transmission Schedule:

All reports by 22:00.

Plans for Tomorrow:

As it's our last day in sim, we'll packed by this afternoon and prepare the Hab for the next rotation.
We've asked Jean Hunter specific guidelines to manage the food left over.

Inventory:

Not applicable

Support Requested:

Not applicable

Miscellaneous:

Commander's Report

Date and time: 

15 December 2011

Written by: 

Charlotte Poupon, Crew 108, Commander


Text:

5 crew members still managing the sim pretty well and are enjoying EVA's
very much, even if the weather is not as bright as at the begining of our rotation.

Somehow, we start to miss Earth, specifically non-dehydrated food !

Tomorrow, we'll get the hab ready to welcome the next crew on Saturday.

Commander waiting for specific guidelines for that…

EVA 22 Report

Date and time: 

15 December 2011

Written by: 

Michael LeClair

Crew Members: 

Michael LeClair (Geologist), Mike Lotto (Engineer)

Time: 10:11 – 12:46

- Depressurization: 10:11 – 10:16

- Turn-around on south end of Radio Ridge road at 12:02

- Travel east on Sagan Street to Factory Bench, arrived 12:19

- Repressurization: 12:41 – 12:46

Site Location: 

Radio Ridge

Transit Mode: 

ATV

Objective: 

Document geology of ridge rock while travelling as far south along Radio Ridge Road as possible.

Results: 

 Site of geological interest located.

Lessons Learned: 

Maintain situational awareness at all times.

EVA 23 Report

Date and time: 

15 December 2011

Written by: 

Michael LeClair

Crew Members: 

Michael LeClair (Geologist), Usha Lingappa (Astrobiologist), Ashley Dale (Engineer)

Time: 

13:45 – 15:55

- Depressurization: 13:45 – 13:50

- Park vehicles at approximately 14:50

- Travel by foot down canyon.

- Repressurization: 15:50 – 15:55

Site Location: 

Radio Ridge

Transit Mode:

ATV

Objective:

Travel to the end of Hubble Highway.

Results:

Site of geological interest located.

Lessons Learned:

Singing is fun.

Engineering Report

Date and time:

December 15 2011, 17:00

Written by: 

Mike Lotto, Crew 108, Engineer

Generator/Electricity:

Did you have any electricity outage? (Yes/No): No

Fuel Status:


Diesel (Full/0.75/0.5/0.25/Empty): 0.25
Propane Tank (Full/0.75/0.5/0.25/Empty): 0.75
Gasoline Tank (Full/0.75/0.5/0.25/Empty): 0.5
Notes/Comments: To prepare for Crew 109, more gasoline for the ATVs is requested.

Oil:


ATV Oil Quantity (# of Quarts in storage): 1.25
Car Oil Quantity (# of Quarts in storage): 1.25

Water Status:


Outside Potable Water Tank Level (inches from bottom): 13
Trailer Potable Water Tank Level (inches from bottom): Approximately 30 inches (or 70% full)
Hab Potable Water Tank Level (inches from bottom): 10
Potable Water Meter Reading: Unknown
In to GreenHab Meter Reading: Offline
Notes/Comments:

GreenHab:


Condition of plants in Tank 1: Offline
Condition of plants in Tank 2: Offline
Condition of plants in Tank 3: Offline
Greenhab Notes/Comments: GreenHab is offline this season.

Observatory:


Telescope Used? (Yes/No): No
Observatory Notes/Comments:

Transportation:


Hab Car (NOT RENTAL)
Used (yes/no): No
Oil Checked (yes/no): No
Tire Status: Good
Odometer at end of day: 125464 miles
Notes/Comments on Hab Car:

ATV 1 (Opportunity):
Used: Yes
Oil Checked: Yes
Fuel Consumed: 1.6 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 12:44
Comments:

ATV 2 (Spirit):
Used: Yes
Oil Checked: Yes
Fuel Consumed: 1 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 08:13
Comments:

ATV 3 (Viking I):
Used: Yes
Oil Checked: Yes
Fuel Consumed: 1.6 gallons
Tire Status: Good
No. of hours ridden since last oil change (HH:MM): 09:29
Comments:

Heating and Ventilation:


Thermostat setting upstairs (Farenheit): 69
Actual temperature upstairs (Farenheit): 66

Thermostat setting downstairs (Farenheit): 55
Actual temperature downstairs (Farenheit): 54

Computers/Networking Infrastructure:


Notes/comments on internet/computers?: None

General Engineering Notes/Comments:


EVA Suit 4 is disabled at the moment; however, it is currently being repaired. A new pack is expected to arrive as well.

Journalist Report

Date and Time: 

15 December 2011, 18:39

Written by: 

Usha Lingappa, Crew 108, Astrobiologist


This was the first morning that we weren't up before the sunrise. Both of today's EVAs took us exploring farther out than we have been before. In the morning Lotto and LeClair went east and in the afternoon Dale, LeClair, and myself went north. LeClair is a champion; he did 6 hours of EVA today. We have now traveled all of the routes documented by previous crews. The engineers have been working on repairing suit 4, and I am wrapping up my work in the lab. We are all sad that the mission is almost over.


Chef's Report

Date:

15 December 2011

Written by:

Charlotte Poupon and Mike Lotto, Crew 108.

1a) Today was a non-cooking day.

2) List the foods served at today's main meal (usually the evening
meal), giving full names of each.

1/2 can of Long Grain & Wild Rice Pilaf

3) List any main meal foods not finished by the crew

none

4) Recipes for anything more complicated than following package
directions exactly, or rehydrating.

none

5) Lessons learned (if any)

6) Comments/questions for Kim and Jean, the food study investigators